OpenStudy (flvs.net):

If 2 over 7 of a ton of concrete covers 7 over 8 of a wall, how many tons of concrete are required to cover the entire wall? A. 1/4 ton B. 16/49 ton C. 5/7 ton D. 65/56 tons

OpenStudy (flvs.net):

@jcoury

OpenStudy (danjs):

\[\frac{ 1~wall }{ 1 }*\frac{ 2/7 ~ ton }{ 7/8~wall } = ~~~ton\]

OpenStudy (anonymous):

cross multiply

OpenStudy (danjs):

no, just calculate

OpenStudy (flvs.net):

How?

OpenStudy (danjs):

oh, to divide a fraction by a fraction, flip the bottom one over and multiply them

OpenStudy (danjs):

\[\frac{ 2/7 }{ 7/8 } = \frac{ 2 }{ 7 }*\frac{ 8 }{ 7 }\]

OpenStudy (flvs.net):

So what do I do?

OpenStudy (danjs):

multiply \[\frac{ 2/7 }{ 7/8 } = \frac{ 2 }{ 7 }*\frac{ 8 }{ 7 }=\frac{ 8*2 }{ 7*7 } = \]

OpenStudy (flvs.net):

Huh?

OpenStudy (danjs):

Dividing Fractions - \[\frac{ \frac{ a }{ b } }{ \frac{ c }{ d } } = \frac{ a }{ b }*\frac{ d }{ c }\]

OpenStudy (danjs):

multiplying fractions - \[\frac{ a }{ b }*\frac{ c }{ d } = \frac{ a*c }{ b*d }\]

OpenStudy (danjs):

so your final answer is (8*2) / (7*7) tons